OpenStudy (anonymous):

Graph the function. Identify the vertex and axis of symmetry. f(x) = –x2 – x + 2 a. axis of symmetry: x = –0.5 vertex: (–0.5, 2.25) b. axis of symmetry: x = 0.5 vertex: (0.5, 2.25) c. axis of symmetry: x = –0.5 vertex: (–0.5, –2.25) d. axis of symmetry: x = 0.5 vertex: (0.5, –2.25)

OpenStudy (raden):

the axis of symmetry = -b/2a from ur equation f(x) = –x2 – x + 2 here, a=-1 and b=-1

OpenStudy (raden):

then to get the vertex, it can be (-b/2a, f(-b/2a))

OpenStudy (anonymous):

thank you